### v1

# A Family of Berndt-Type Integrals and Associated Barnes Multiple Zeta Functions

Abstract. In this paper, we focus on calculating a specific class of Berndt integrals, which exclusively involves (hyperbolic) cosine functions. Initially, this integral is transformed into a Ramanujan-type hyperbolic (infinite) sum via contour integration. Subsequently, a function incorporating $\theta$ is defined. By employing the residue theorem, the mixed Ramanujan-type hyperbolic (infinite) sum with both hyperbolic cosine and hyperbolic sine in the denominator is converted into a simpler Ramanujan-type hyperbolic (infinite) sum, which contains only hyperbolic cosine or hyperbolic sine in the denominator. The simpler Ramanujan-type hyperbolic (infinite) sum is then evaluated using Jacobi elliptic functions, Fourier series expansions, and Maclaurin series expansions. Ultimately, the result is expressed as a rational polynomial of $\Gamma(1/4)$ and $\pi^{-1/2}$. Additionally, the integral is related to the Barnes multiple zeta function, which provides an alternative method for its calculation.
Keywords: Berndt-type integral, Hyperbolic functions, Contour integration, Jacobi elliptic functions, Barnes multiple zeta functions.
AMS Subject Classifications (2020): 33E05, 33E20, 44A05, 11M99.
1 Introduction
In [22], the second author of this paper and Zhao referred to integrals of the following form as $m$-th order Berndt-type integrals, since it was approximately a decade ago that Professor Berndt first systematically studied the case when $m=1$. For positive integers $s$ and $m$, the Berndt-type integrals of order $m$ is defined by
${\rm BI}^{\pm}(s,m):=\displaystyle\int_{0}^{\infty}\displaystyle\frac{x^{s-1}\mathrm{d}x}{(\cos x\pm\cosh x)^{m}},$ (1.1)
where $s\geq 1$ and $m\geq 1$ for the “$+$’ case and $s\geq 2m+1$ otherwise. In fact, the study of Berndt-type integrals can be traced back to 1916, when the renowned Indian mathematician Ramanujan investigated the following improper integrals involving trigonometric and hyperbolic functions. He conjectured the integral identity (see [14, pp. 325-326])
$\int_{0}^{\infty}\displaystyle\frac{\sin(nx)}{x(\cos x+\cosh x)}\mathrm{d}x=\displaystyle\frac{\pi}{4}\quad(n\in\mathbb{N}),$
which was only rigorously proven four years later by Wilkinson [19].
About 80 years afterward, the study of such integrals gained renewed interest with the work of Ismail and Valent [10], who discovered another remarkable integral (Ismail-type integrals), exemplified by
$\displaystyle\int_{-\infty}^{\infty}{\displaystyle\frac{\mathrm{d}t}{\cos\left(K\sqrt{t}\right)+\cosh\left(K^{\prime}\sqrt{t}\right)}=2}.$ (1.2)
Here $K\equiv K(x)$ denotes the complete elliptic integral of the first kind and $K^{\prime}=K(1-x)$. Kuznetsov [12] generalized Ismail-type integrals using contour integration and theta functions, proving
$\displaystyle\frac{1}{2}\displaystyle\int_{-\infty}^{\infty}\displaystyle\frac{t^{n}\mathrm{d}t}{\cos(K\sqrt{t})+\cosh(K^{\prime}\sqrt{t})}=(-1)^{n}\displaystyle\frac{d^{2n+1}}{du^{2n+1}}\displaystyle\frac{{\rm sn}(u,k)}{{\rm cd}(u,k)}|_{u=0},$ (1.3)
where $x=k^{2}\ (0<k<1)$ and ${\rm sn}(u,k)$ is the Jacobi elliptic function defined through the inversion of the elliptic integral
$u=\displaystyle\int_{0}^{\varphi}\displaystyle\frac{\mathrm{d}t}{\sqrt{1-k^{2}\sin^{2}t}}\quad(0<k^{2}<1),$ (1.4)
that is, ${\rm sn}(u):=\sin\varphi$. As before, $k$ is referred to as the elliptic modulus. We also write $\varphi={\rm am}(u,k)={\rm am}(u)$ and call it the Jacobi amplitude. The Jacobi elliptic function ${\rm cd}(u,k)$ can be defined as follows
${\rm cd}(u,k):=\displaystyle\frac{{\rm cn}(u,k)}{{\rm dn}(u,k)},$
where ${\rm cn}(u,k):=\sqrt{1-{\rm sn}^{2}(u,k)}$ and ${\rm dn}(u,k):=\sqrt{1-k^{2}{\rm sn}^{2}(u,k)}$.
In the past decade, research on these analogous integrals has gradually attracted increasing interest, yielding notable progress. Recent advances in this direction can be found in the works of [7, 13, 16, 17, 21, 22, 24] and the references therein. A significant turning point in this line of research occurred in 2016 when Professor Berndt [6] developed a systematic approach to study ${\rm BI}^{\pm}(s,1)$. Berndt further advanced the field by evaluating such integrals directly using Cauchy’s residue theory, Fourier series expansions, and Maclaurin series expansions of Jacobi elliptic functions. In their series of works [17, 21, 22], Rui, Xu, and Zhao developed an extended framework based on Berndt’s approach, which led to fundamental structural theorems for higher-order Berndt-type integrals. Letting $X=\Gamma^{4}(1/4)$ and $Y=\pi^{-1}$, they proved that for all integers $m\geq 1$ and $p\geq[m/2]$ (see [22, Thms 1.1 and 1.2]),
${\rm BI}^{+}(4p+2,m)\in\mathbb{Q}[X,Y]\quad\text{and}\quad{\rm BI}^{-}(a,m)\in\mathbb{Q}[X,Y]\quad(0<a-2m\equiv 1\pmod{4}),$
and the optimal range for the degrees of the polynomials in $X$ and $Y$ has been precisely determined.
Very recently, a groundbreaking connection between Berndt-type integrals and Barnes multiple zeta functions (See Section 7) was established by Bradshaw and Vignat [7, Cor. 5], who also linked Jacobi elliptic functions to Ismail-type integrals: for $n\in\mathbb{N}_{0}$,
$\displaystyle\frac{1}{2}\displaystyle\int_{-\infty}^{\infty}\displaystyle\frac{x^{n+1}dx}{\cos(\sqrt{x}K)-\cosh(\sqrt{x}K^{\prime})}=(-1)^{n+1}4\displaystyle\frac{d^{2n+1}}{du^{2n+1}}\displaystyle\frac{{\rm sn}^{2}(u,k)}{{\rm cd}^{2}(u,k){\rm sd}(2u,k)}|_{u=0},$ (1.5)
where the Jacobi elliptic function ${\rm sd}(u)\equiv{\rm sd}(u,k)$ is defined by
${\rm sd}(u,k):=\displaystyle\frac{{\rm sn}(u,k)}{{\rm dn}(u,k)}.$ (1.6)
As shown in the preceding analysis, the integrands of the previously studied Berndt-type integrals consist solely of additive or subtractive combinations of a single trigonometric cosine and a single hyperbolic cosine function
In this paper, we focus on integrals of the form:
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{s}\mathrm{d}x}{\left[\cosh(2x)-\cos(2x)\right]\left[\cosh x+\cos x\right]}},$ (1.7)
which feature purely (hyperbolic) cosine integrands and the integrand involves the product of two terms: (i) the sum of a cosine and a hyperbolic cosine function, and (ii) their difference. By leveraging contour integration and series expansions of Jacobi elliptic functions, we derive structural theorems and explicit evaluations for these Berndt-type integrals. We are going to prove the following evaluations (see Theorem 6.1): let $\Gamma:=\Gamma(1/4)$, for $m\in\mathbb{N}$,
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m-1}\mathrm{d}x}{\left[\cosh(2x)-\cos(2x)\right]\left[\cosh x+\cos x\right]}}$ $\in\mathbb{Q}\displaystyle\frac{\Gamma^{8m-4}}{\pi^{2m-1}}+\displaystyle\frac{\mathbb{Q}}{\sqrt{2}}\displaystyle\frac{\Gamma^{8m-2}}{\pi^{(4m-1)/2}}$
$\quad+\mathbb{Q}\displaystyle\frac{\Gamma^{8m}}{\pi^{2m}}+\displaystyle\frac{\mathbb{Q}}{\sqrt{2}}\displaystyle\frac{\Gamma^{8m+2}}{\pi^{(4m+3)/2}}+\mathbb{Q}\displaystyle\frac{\Gamma^{8m+4}}{\pi^{2m+3}}.$ (1.8)
Furthermore, we put forward the specific coefficients preceding each term and set up some explicit relations between the Berndt-type integrals and Barnes multiple zeta functions. Additionally, we establish their relationship with Barnes multiple zeta functions, offering new closed-form evaluations and insights (see Theorem 7.2).
2 Some Preliminary Results
In this section, we present fundamental mathematical tools and special functions that form the basis for our subsequent analysis. Let ${}_{2}F_{1}\left(a,b;c;x\right)$ denote the Gaussian hypergeometric function defined by
${{}_{2}}F_{1}(a,b;c;x)=\displaystyle\sum\limits_{n=0}^{\infty}\displaystyle\frac{(a)_{n}(b)_{n}}{(c)_{n}}\displaystyle\frac{x^{n}}{n!}\quad(a,b,c\in\mathbb{C}),$
where $\left(a\right)_{0}:=1$ and $\left(a\right)_{n}=\displaystyle\frac{\Gamma\left(a+n\right)}{\Gamma\left(a\right)}=\prod_{j=0}^{n-1}{\left(a+j\right)}\quad(n\in\mathbb{N}),$ and $\Gamma(x)$ is the Gamma function defined by
$\Gamma\left(z\right):=\displaystyle\int_{0}^{\infty}{e^{-t}t^{z-1}\mathrm{d}t}\quad\left(\Re\left(z\right)>0\right).$
We know that ${}_{2}F_{1}\left(a,b;c;x\right)$ plays a crucial role in evaluating elliptic integrals, particularly the complete elliptic integral of the first and second kinds are defined by (Whittaker and Watson [18])
$K:=K\left(k^{2}\right)=\displaystyle\int_{0}^{\pi/2}{\displaystyle\frac{\mathrm{d}\varphi}{\sqrt{1-k^{2}\sin^{2}\varphi}}=\displaystyle\frac{\pi}{2}}\,\,_{2}F_{1}\left(\displaystyle\frac{1}{2},\displaystyle\frac{1}{2};1;k^{2}\right),$
$E:=E(k^{2}):=\displaystyle\int_{0}^{\pi/2}\sqrt{1-k^{2}\sin^{2}\varphi}\mathrm{d}\varphi=\displaystyle\frac{\pi}{2}{{}_{2}}F_{1}\left(-\displaystyle\frac{1}{2},\displaystyle\frac{1}{2};1;k^{2}\right).$
Following Ramanujan’s notation, we define key variables and their relationships:
$x:=k^{2},\ y\left(x\right):=\pi K^{\prime}/K,\ q\equiv q\left(x\right):=e^{-y},\ z:=z\left(x\right)=2K/\pi,\ z^{\prime}=dz/dx,$ (2.9)
where $K^{\prime}=K(1-x)$. Then
$\frac{\mathrm{d}^{n}z}{\mathrm{d}x^{n}}=\displaystyle\frac{\left(1/2\right)_{n}^{2}}{n!}\,_{2}F_{1}\left(\displaystyle\frac{1}{2}+n,\displaystyle\frac{1}{2}+n,1+n;x\right).$
Applying the identity ([2])
$\,{}_{2}F_{1}\left(a,b,\displaystyle\frac{a+b+1}{2};\displaystyle\frac{1}{2}\right)=\displaystyle\frac{\Gamma\left(\displaystyle\frac{1}{2}\right)\Gamma\left(\displaystyle\frac{a+b+1}{2}\right)}{\Gamma\left(\displaystyle\frac{a+1}{2}\right)\Gamma\left(\displaystyle\frac{b+1}{2}\right)},$
these satisfy important differential properties
$\displaystyle\frac{d^{n}z}{dx^{n}}\mid_{x=\tfrac{1}{2}}=\displaystyle\frac{\left(1/2\right)_{n}^{2}\sqrt{\pi}}{\Gamma^{2}\left(\displaystyle\frac{n}{2}+\displaystyle\frac{3}{4}\right)}.$ (2.10)
In particular, taking $x=\displaystyle\frac{1}{2}$ and $n=0,1,2$ in (2.10), we obtain that
$y=\pi,\quad z\left(\displaystyle\frac{1}{2}\right)=\displaystyle\frac{\Gamma^{2}(1/4)}{2\pi^{3/2}},\quad z^{\prime}\left(\displaystyle\frac{1}{2}\right)=\displaystyle\frac{4\sqrt{\pi}}{\Gamma^{2}(1/4)},\quad z^{\prime\prime}\left(\displaystyle\frac{1}{2}\right)=\displaystyle\frac{\Gamma^{2}(1/4)}{2\pi^{3/2}},$
which are essential for our explicit evaluations. The derivative relation:
$\frac{dx}{dy}=-x(1-x)z^{2}$ (2.11)
provides a crucial connection between these variables that we exploit in later calculations.
3 Some Definitions and Lemmas
This section establishes the core mathematical framework for our subsequent analysis through carefully constructed definitions and technical lemmas.
Definition 3.1.
Let $s\in\mathbb{C}$ and $a,b,\theta\in\mathbb{R}$ with $|\theta|<2b\pi$ and $a,b\neq 0$. Define
$Z_{1}(s,\theta;a,b)$ $:=\displaystyle\frac{\pi^{4}\sinh(\theta s)}{\sin(a\pi s)\sinh(b\pi s)\cosh^{2}(b\pi s)},$ $Z_{2}(s,\theta;a,b):$ $=\displaystyle\frac{\pi^{4}\cosh(\theta s)}{\cos(a\pi s)\sinh(b\pi s)\cosh^{2}(b\pi s)},$
$Z_{3}(s,\theta;a,b)$ $:=\displaystyle\frac{\pi^{4}\sinh(\theta s)}{\cos(a\pi s)\sinh^{2}(b\pi s)\cosh(b\pi s)},$ $Z_{4}(s,\theta;a,b):$ $=\displaystyle\frac{\pi^{4}\sinh(\theta s)}{\cos(a\pi s)\cosh^{3}(b\pi s)}.$
Definition 3.2.
Let $m\in\mathbb{N}$ and $p\in\mathbb{Z}$. Define
$G_{p,m}(y)$ $:=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}n^{p}}{\sinh(ny)\cosh^{m}(ny)},\quad X_{p,m}(y):=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}n^{p}}{\sinh^{m}(ny)},$
$G_{p,m}^{\prime}(y)$ $:=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{p-1}}{\sinh((2n-1)y/2)\cosh^{m}((2n-1)y/2)},$
$\overline{G}_{p,m}(y)$ $:=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{p}}{\sinh^{m}((2n-1)y/2)\cosh((2n-1)y/2)},$
$Y_{p,m}(y)$ $:=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{p-1}}{\sinh^{m}((2n-1)y/2)},$
$\mathrm{DY}_{p,m}(y)$ $:=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{p}\cosh((2n-1)y/2)}{\sinh^{m}((2n-1)y/2)},$
$X_{p,m}^{\prime}(y)$ $:=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{p}}{\cosh^{m}((2n-1)y/2)},$
$\mathrm{DX}_{p,m}^{\prime}(y)$ $:=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{p+1}\sinh((2n-1)y/2)}{\cosh^{m}((2n-1)y/2)},$
$B_{p,m}(y)$ $:=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}n^{p-1}}{\cosh^{m}(ny)},\quad\mathrm{DB}_{p,m}(y):=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}n^{p}\sinh(ny)}{\cosh^{m}{ny}}.$
Lemma 3.1.
(Residue Theorem [9]) Let $\xi(s)$ be a kernel function and let $r(s)$ be a function which is $O(s^{-2})$ at infinity. Then
$\sum_{\begin{subarray}{c}\alpha\in O\\ \end{subarray}}\text{Res}\left(r(s)\xi(s),s=\alpha\right)+\displaystyle\sum_{\begin{subarray}{c}\beta\in S\\ \end{subarray}}\text{Res}\left(r(s)\xi(s),s=\beta\right)=0,$ (3.12)
where $S$ is the set of poles of $r(s)$ and $O$ is the set of poles of $\xi(s)$ that are not poles $r(s)$. Here $\text{Res}\left(r(s),s=\alpha\right)$ denotes the residue of $r(s)$ at $s=\alpha$. The kernel function $\xi(s)$ is meromorphic in the whole complex plane and satisfies $\xi(s)=o(s)$ over an infinite collection of circles $|s|=\rho_{k}$ with $\rho_{k}\rightarrow\infty$.
Lemma 3.2.
([23, Thm. 2.4]) Let $x,y,z$ and $z^{\prime}$ satisfy (2.9). Given the formula $\Omega(x,e^{-y},z,z^{\prime})=0$, we have the transformation formula
$\Omega\left(1-x,e^{-\pi^{2}/y},yz/\pi,\displaystyle\frac{1}{\pi}\left(\displaystyle\frac{1}{x\left(1-x\right)z}-yz^{\prime}\right)\right)=0.$ (3.13)
Lemma 3.3.
(cf.[20, Lem. 1.2]) Let $n$ be an integer, then the following formulas hold:
$\displaystyle\frac{\pi}{\sin(\pi s)})\overset{s\rightarrow n}{=}(-1)^{n}\left(\displaystyle\frac{1}{s-n}+2\displaystyle\sum_{k=1}^{\infty}\overline{\zeta}(2k)(s-n)^{2k-1}\right),$ (3.14)
$\displaystyle\frac{\pi}{\sinh(\pi s)}\overset{s\rightarrow ni}{=}(-1)^{n}\left(\displaystyle\frac{1}{s-ni}+2\displaystyle\sum_{k=1}^{\infty}(-1)^{k}\overline{\zeta}(2k)(s-ni)^{2k-1}\right),$ (3.15)
$\displaystyle\frac{\pi}{\cos(\pi s)}\overset{s\rightarrow n-1/2}{=}(-1)^{n}\left\{\displaystyle\frac{1}{s-\displaystyle\frac{2n-1}{2}}+2\displaystyle\sum_{k=1}^{\infty}\overline{\zeta}(2k)\left(s-\displaystyle\frac{2n-1}{2}\right)^{2k-1}\right\},$ (3.16)
$\displaystyle\frac{\pi}{\cosh(\pi s)}\overset{s\rightarrow(n-1/2)i}{=}(-1)^{n}i\left\{\displaystyle\frac{1}{s-\displaystyle\frac{2n-1}{2}i}+2\displaystyle\sum_{k=1}^{\infty}(-1)^{k}\overline{\zeta}(2k)\left(s-\displaystyle\frac{2n-1}{2}i\right)^{2k-1}\right\}.$ (3.17)
Lemma 3.4.
([21, Eqs. (4.4)-(4.7)]) Let $n\in\mathbb{Z}$. We have
$\displaystyle\frac{\left(-1\right)^{n}}{\sin\left(\displaystyle\frac{1\pm i}{2}z\right)}\overset{z\rightarrow\left(1\mp i\right)n\pi}{=}\displaystyle\frac{1\mp i}{z-\left(1\mp i\right)n\pi}+2\displaystyle\sum_{k=1}^{\infty}{\displaystyle\frac{\overline{\zeta}\left(2k\right)}{\pi^{2k}}\left(\displaystyle\frac{1\pm i}{2}\right)^{2k-1}\left(z-\left(1\mp i\right)n\pi\right)^{2k-1}},$ (3.18)
$\displaystyle\frac{\left(-1\right)^{n-1}}{\cos\left(\displaystyle\frac{1\pm i}{2}z\right)}\overset{z\rightarrow\left(1\mp i\right)\tilde{n}\pi}{=}\displaystyle\frac{1\mp i}{z-\left(1\mp i\right)\tilde{n}\pi}-2\displaystyle\sum_{k=1}^{\infty}{\displaystyle\frac{\overline{\zeta}\left(2k\right)}{\pi^{2k}}\left(\displaystyle\frac{1\pm i}{2}\right)^{2k-1}\left(z-\left(1\mp i\right)\tilde{n}\pi\right)^{2k-1}},$ (3.19)
where $\tilde{n}:=n-\displaystyle\frac{1}{2}$. Here $\zeta(s)$ and $\overline{\zeta}(s)$ stand for the Riemann zeta function and alternating Riemann zeta function, respectively, which are defined by
$\zeta\left(s\right):=\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{1}{n^{s}}}\quad(\Re(s)>1)\quad\text{and}\quad\bar{\zeta}\left(s\right):=\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{\left(-1\right)^{n-1}}{n^{s}}}\quad\left(\Re\left(s\right)>0\right).$
When $s=2m$ is an even, Euler proved the famous formula
$\zeta\left(2m\right)=\displaystyle\frac{\left(-1\right)^{m-1}B_{2m}\left(2\pi\right)^{2m}}{2\left(2m\right)!},$
where $B_{2m}\in\mathbb{Q}$ are Bernoulli numbers defined by the generating function
$\frac{x}{e^{x}-1}=\displaystyle\sum_{n=0}^{\infty}{\displaystyle\frac{B_{n}}{n!}x^{n}.}$
Lemma 3.5.
([8]) The Maclaurin series of ${\rm cd}(u)$ and ${\rm nd}(u)$ have the forms
$\mathrm{cd}(u)=\displaystyle\sum_{n=0}^{\infty}{\mathrm{S}_{2n}\left(x\right)\displaystyle\frac{\left(-1\right)^{n}u^{2n}}{\left(2n\right)!}}\quad\text{and}\quad\mathrm{nd}(u)=\displaystyle\sum_{n=0}^{\infty}{\mathrm{A}_{2n}\left(x\right)\displaystyle\frac{\left(-1\right)^{n}u^{2n}}{\left(2n\right)!}},$ (3.20)
where $\mathrm{nd}(u):=1/\mathrm{dn}(u)$, $\mathrm{S}_{2n}\left(x\right),\mathrm{A}_{2n}\left(x\right)\in\mathbb{Z}[x].$
4 Berndt-Type Integrals via Hyperbolic Series
In this section, we establish exact correspondences between Berndt-type integrals and Ramanujan-type hyperbolic series through the contour integration techniques. Our approach reveals deep connections between these apparently distinct mathematical objects.
Theorem 4.1.
For any positive integer $p\geq 3$, the following identity holds:
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{p}\mathrm{d}x}{\left[\cosh(2x)-\cos(2x)\right]\left[\cosh x+\cos x\right]}}+(-i)^{p+1}\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{p}\mathrm{d}x}{\left[\cosh(2x)-\cos(2x)\right]\left[\cosh x+\cos x\right]}}$
$=\displaystyle\frac{1}{2^{p+1}}ip((1-i))^{p-1}\pi^{p}\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{\left(-1\right)^{n}(2n-1)^{p-1}}{\sinh\left((2n-1)\pi/2\right)\cosh^{2}\left((2n-1)\pi/2\right)}}$
$\quad+\displaystyle\frac{1}{8}((1-i))^{p+1}\pi^{p+1}\displaystyle\sum_{m=1}^{\infty}{\displaystyle\frac{\left(-1\right)^{m}m^{p}}{\sinh(m\pi)\cosh^{2}(m\pi)}}-\displaystyle\frac{1}{2^{p+1}}i(1-i)^{p-1}\pi^{p+1}\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{\left(-1\right)^{n}(2n-1)^{p}}{\cosh^{3}\left((2n-1)\pi/2\right)}}$
$\quad-\displaystyle\frac{1}{2^{p+2}}i((1-i))^{p-1}\pi^{p+1}\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{\left(-1\right)^{n}(2n-1)^{p}}{\sinh^{2}\left((2n-1)\pi/2\right)\cosh\left((2n-1)\pi/2\right)}}.$ (4.21)
Proof.
Let $z=x+iy$ for $x,y\in\mathbb{R}$. Consider the contour integral
$\mathscr{I}_{a}=\displaystyle\lim_{R\rightarrow\infty}\displaystyle\int_{\mathrm{C}_{R}}{\displaystyle\frac{z^{p}\mathrm{d}z}{\left[\cosh(2z)-\cos(2z)\right]\left[\cosh z+\cos z\right]}}=\displaystyle\lim_{R\rightarrow\infty}\displaystyle\int_{\mathrm{C}_{R}}{F\left(z\right)}\mathrm{d}z,$ (4.22)
where $C_{R}$ denotes the quarter-circular contour consisting of the interval $[0,R]$, the qiarter-circle $\Gamma_{R}$ with $|z|=R\ (-\pi/2\leq argz\leq 0)$, and the segment $[-iR,0]$, the contour is illustrated below:
Obviously, there exist poles in case that
$\left[\cosh(2z)-\cos(2z)\right]\left[\cosh z+\cos z\right]=16\sin\left\{\displaystyle\frac{1+i}{2}z\right\}\sin\left\{\displaystyle\frac{1-i}{2}z\right\}\cos^{2}\left\{\displaystyle\frac{1+i}{2}z\right\}\cos^{2}\left\{\displaystyle\frac{1-i}{2}z\right\}=0.$
The poles enclosed by $C_{R}$ are located at $z_{m}=m\pi(1-i)\ (m\geq 1,\ |z_{m}|<R)$ and $z_{n}=(2n-1)\pi(1-i)/2\ (n\geq 1,\ |z_{n}|<R)$. Using Lemma 3.4, the residues of $\text{Res}[F(z),z]$ at these poles are given by:
$\mathrm{Res}\left[F\left(z\right),z_{m}\right]$ $=\displaystyle\frac{\left(-1\right)^{m}(1-i)\left[(1-i)m\pi\right]^{p}}{16\sin\left(-im\pi\right)\cos^{2}\left(m\pi\right)\cos^{2}\left(-im\pi\right)}=\displaystyle\frac{\left(-1\right)^{m-1}(1-i)\left[(1-i)m\pi\right]^{p}}{16i\sinh\left(m\pi\right)\cosh^{2}\left(m\pi\right)},$ (4.23)
$\mathrm{Res}\left[F\left(z\right),z_{n}\right]$ $=(1-i)^{2}\left[\displaystyle\frac{p\left[(2n-1)\pi(1-i)/2\right]^{p-1}}{16(-1)^{n-1}\left[-i\sinh((2n-1)\pi/2)\right]\cosh^{2}((2n-1)\pi/2)}\right]$
$\quad-(1-i)^{2}\left[\displaystyle\frac{(1-i)\left[(2n-1)\pi(1-i)/2\right]^{p}}{32(-1)^{n-1}\left[-i\sinh((2n-1)\pi/2)\right]^{2}\cosh((2n-1)\pi/2)}\right]$
$\quad+(1-i)^{2}\left[\displaystyle\frac{(1-i)\left[(2n-1)\pi(1-i)/2\right]^{p}}{16(-1)^{n-1}\cosh^{3}((2n-1)\pi/2)}\right].$ (4.24)
As $R\rightarrow\infty$, we get
$\int_{\Gamma_{R}}{\displaystyle\frac{z^{p}\mathrm{d}z}{\left[\cosh(2z)-\cos(2z)\right]\left[\cosh z+\cos z\right]}}=o\left(1\right).$
Applying the residue theorem and taking the limit $R\rightarrow\infty$, we obtain
$-2\pi i\displaystyle\sum_{m=1}^{\infty}{\mathrm{Res}\left[F\left(z\right),z_{m}\right]}-2\pi i\displaystyle\sum_{n=1}^{\infty}{\mathrm{Res}\left[F\left(z\right),z_{n}\right]}=\displaystyle\lim_{R\rightarrow\infty}\displaystyle\int_{\mathrm{C}_{R}}{\displaystyle\frac{z^{p}\mathrm{d}z}{\left[\cosh(2z)-\cos(2z)\right]\left[\cosh z+\cos z\right]}}$
$=\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{p}\mathrm{d}x}{\left[\cosh(2x)-\cos(2x)\right]\left[\cosh x+\cos x\right]}}+i\displaystyle\int_{0}^{\infty}{\displaystyle\frac{\left(-ix\right)^{p}\mathrm{d}x}{\left[\cosh(-2ix)-\cos(-2ix)\right]\left[\cosh(-ix)+\cos(-ix)\right]}}$
$=\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{p}\mathrm{d}x}{\left[\cosh(2x)-\cos(2x)\right]\left[\cosh x+\cos x\right]}}+(-i)^{p+1}\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{p}\mathrm{d}x}{\left[\cosh(2x)-\cos(2x)\right]\left[\cosh(x)+\cos(x)\right]}}.$
Combining these results yield the desired identity. ∎
Theorem 4.2.
For $a,b,\theta\in\mathbb{R}$ and $|\theta|<2b\pi$, $a,b\neq 0$,
$b^{2}\pi\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}\sinh(n\theta/a)}{\sinh(bn\pi/a)\cosh^{2}(bn\pi/a)}+ab\pi\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}\sin(n\theta/b)}{\sinh(an\pi/b)}-a\theta\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}\cos((2n-1)\theta/(2b))}{\sinh((2n-1)a\pi/(2b))}$
$+a^{2}\pi\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}\sin((2n-1)\theta/(2b))\cosh((2n-1)a\pi/(2b))}{\sinh^{2}((2n-1)a\pi/(2b))}+\displaystyle\frac{\theta b}{2}=0,$ (4.25)
$b^{2}\pi\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}\cosh((2n-1)\theta/(2a))}{\sinh((2n-1)b\pi/(2a))\cosh^{2}((2n-1)b\pi/(2a))}+a\theta\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}\sin((2n-1)\theta/(2b))}{\cosh((2n-1)a\pi/(2b))}$
$+a^{2}\pi\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}\cos((2n-1)\theta/(2b))\sinh((2n-1)a\pi/(2b))}{\cosh^{2}((2n-1)a\pi/(2b))}+ab\pi\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}\cos(n\theta/b)}{\cosh(an\pi/b)}$
$\quad+\displaystyle\frac{ab\pi}{2}=0,$ (4.26)
$b^{2}\pi\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}\sinh((2n-1)\theta/(2a))}{\sinh^{2}((2n-1)b\pi/(2a))\cosh((2n-1)b\pi/(2a))}+a\theta\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}\cos(n\theta/b)}{\cosh(an\pi/b)}$
$-a^{2}\pi\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}\sin(n\theta/b)\sinh(an\pi/b)}{\cosh^{2}(an\pi/b)}+ab\pi\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}\sin((2n-1)\theta/(2b))}{\cosh((2n-1)a\pi/(2b))}+\displaystyle\frac{a\theta}{2}=0.$ (4.27)
Proof.
This proof relies on analyzing the functions $Z_{i}(s,\theta;a,b)$($j=1,2,3,4$) and the usual residue computation. These functions are meromorphic in the entire complex plane with some simple poles at specific points.
First, we note that function $Z_{1}(s,\theta;a,b)$ has poles at $s=\pm\displaystyle\frac{n}{a}$ (simple poles), $s=\pm\displaystyle\frac{ni}{b}$ (simple poles), $s=\pm\displaystyle\frac{(2n-1)i}{2b}$ (double poles), $s=0$ (simple poles), where $n\in\mathbb{N}$.
As $s=\pm\displaystyle\frac{n}{a}$ and by using Lemma 3.3, the residue is
$\mathrm{Res}\left[Z_{1},s=\pm\displaystyle\frac{n}{a}\right]=\displaystyle\frac{\pi^{3}}{a}\displaystyle\frac{(-1)^{n}\sinh(n\theta/a)}{\sinh(bn\pi/a)\cosh^{2}(bn\pi/a)}.$ (4.28)
As $s=\pm\displaystyle\frac{ni}{b}$ and the residue is
$\mathrm{Res}\left[Z_{1},s=\pm\displaystyle\frac{ni}{b}\right]=\displaystyle\frac{\pi^{3}}{b}\displaystyle\frac{(-1)^{n}\sin(n\theta/a)}{\sinh(an\pi/b)}.$ (4.29)
From Lemma 3.3, for all $n\in\mathbb{Z}$ we may deduce the asymptotic expansions of a few reciprocal quadratic trigonometric and hyperbolic functions as follows
$\left(\displaystyle\frac{\pi}{\cosh(b\pi s)}\right)^{2}$ $=-\displaystyle\frac{1}{\left(bs-\displaystyle\frac{2n-1}{2}i\right)^{2}}+2\zeta(2)-6\zeta(4)\left(bs-\displaystyle\frac{2n-1}{2}i\right)^{2}$
$\quad+o\left(\left(bs-\displaystyle\frac{2n-1}{2}i\right)^{2}\right)\quad\text{as }s\rightarrow\displaystyle\frac{2n-1}{2b}i.$ (4.30)
As $s=\pm\displaystyle\frac{(2n-1)i}{2b}$ and the residue is
$\mathrm{Res}\left[Z_{1},s=\pm\displaystyle\frac{(2n-1)i}{2b}\right]$ $=-\displaystyle\frac{\pi^{2}\theta}{b^{2}}\displaystyle\frac{(-1)^{n}\cos((2n-1)\theta/(2b))}{\sinh((2n-1)a\pi/(2b))}$
$\quad+\displaystyle\frac{a\pi^{3}}{b^{2}}\displaystyle\frac{(-1)^{n}\sin((2n-1)\theta/(2b))\cosh((2n-1)a\pi/(2b))}{\sinh^{2}((2n-1)a\pi/(2b))}.$ (4.31)
Furthermore, applying (3.14) and (3.15) we deduce the expansion
$Z_{1}(s,\theta;a,b)=\displaystyle\frac{\pi^{2}}{\cosh^{2}(b\pi s)}\left(\displaystyle\frac{\theta}{ab}\displaystyle\frac{1}{s}+\displaystyle\frac{\theta^{3}}{6ab}s-\left(\displaystyle\frac{b}{a}-\displaystyle\frac{a}{b}\right)\zeta(2)\theta s+o(1)\right),\quad s\to 0.$
Therefore, as $s=0$ and the residue is
$\mathrm{Res}\left[Z_{1},s=0\right]=\displaystyle\frac{\theta\pi^{2}}{ab}.$ (4.32)
By Lemma 3.1, summing the five contributions (4.28)-(4.32) yields the desired result (4.25).
Secondly, The function $Z_{2}(s,\theta;a,b)$ possesses poles at $s=\pm\displaystyle\frac{2n-1}{2a}$ (simple poles), $s=\pm\displaystyle\frac{ni}{b}$ (simple poles), $s=\pm\displaystyle\frac{(2n-1)i}{2b}$ (double poles), $s=0$ (simple poles), where $n\in\mathbb{N}$.
As $s=\pm\displaystyle\frac{2n-1}{2a}$ and by applying Lemma 3.3, the residue is
$\mathrm{Res}\left[Z_{2},s=\pm\displaystyle\frac{2n-1}{2a}\right]=\displaystyle\frac{\pi^{3}}{a}\displaystyle\frac{(-1)^{n}\cosh((2n-1)\theta/(2a))}{\sinh((2n-1)b\pi/(2a))\cosh^{2}((2n-1)b\pi/(2a))}.$ (4.33)
As $s=\pm\displaystyle\frac{ni}{b}$ and the residue is
$\mathrm{Res}\left[Z_{2},s=\pm\displaystyle\frac{ni}{b}\right]=\displaystyle\frac{\pi^{3}}{b}\displaystyle\frac{(-1)^{n}\cos(n\theta/b)}{\cosh(an\pi/b)}.$ (4.34)
As $s=\pm\displaystyle\frac{(2n-1)i}{2b}$, for this double pole, we employ the expansion from (4) to obtain
$\mathrm{Res}\left[Z_{2},s=\pm\displaystyle\frac{(2n-1)i}{2b}\right]$ $=\displaystyle\frac{\pi^{2}\theta}{b^{2}}\displaystyle\frac{(-1)^{n}\sin((2n-1)\theta/(2b))}{\cosh((2n-1)a\pi/(2b))}$
$\quad+\displaystyle\frac{a\pi^{3}}{b^{2}}\displaystyle\frac{(-1)^{n}\cos((2n-1)\theta/(2b))\sinh((2n-1)a\pi/(2b))}{\cosh^{2}((2n-1)a\pi/(2b))}.$ (4.35)
As $s=0$, using the expansions from (3.15) and (3.16), we derive
$Z_{2}(s,\theta;a,b)=\displaystyle\frac{\pi^{3}\cosh(\theta s)}{\cos(a\pi s)\cosh^{2}(b\pi s)}\left(\displaystyle\frac{1}{b}\displaystyle\frac{1}{s}-\zeta(2)bs+o(1)\right),\quad s\to 0,$
which gives the residue
$\mathrm{Res}\left[Z_{2},s=0\right]=\displaystyle\frac{\pi^{3}}{b}.$ (4.36)
By Lemma 3.1, the sum of these four residue contributions (4.33)-(4.36) establishes the desired identity (4.26).
At last, the function $Z_{3}(s,\theta;a,b)$ has poles at $s=\pm\displaystyle\frac{2n-1}{2a}$ (simple poles), $s=\pm\displaystyle\frac{(2n-1)i}{2b}$ (simple poles), $s=\pm\displaystyle\frac{ni}{b}$ (double poles), $s=0$ (simple poles), where $n\in\mathbb{N}$.
As $s=\pm\displaystyle\frac{2n-1}{2a}$ and by using Lemma 3.3, we compute the residue at these simple poles
$\mathrm{Res}\left[Z_{3},s=\pm\displaystyle\frac{2n-1}{2a}\right]=\displaystyle\frac{\pi^{3}}{a}\displaystyle\frac{(-1)^{n}\sinh((2n-1)\theta/(2a))}{\sinh^{2}((2n-1)b\pi/(2a))\cosh((2n-1)b\pi/(2a))}.$ (4.37)
As =±$=\pm\displaystyle\frac{(2n-1)i}{2b}$, the simple poles contribute
$\mathrm{Res}\left[Z_{3},s=\pm\displaystyle\frac{(2n-1)i}{2b}\right]=\displaystyle\frac{\pi^{3}}{b}\displaystyle\frac{(-1)^{n}\sin((2n-1)\theta/(2b))}{\cosh((2n-1)a\pi/(2b))}.$ (4.38)
As $s=\pm\displaystyle\frac{ni}{b}$, these double poles, we apply the asymptotic expansion from Lemma 3.3
$\left(\displaystyle\frac{\pi}{\sinh(b\pi s)}\right)^{2}$ $=\displaystyle\frac{1}{\left(bs-ni\right)^{2}}-2\zeta(2)+6\zeta(4)\left(bs-ni\right)^{2}+o\left(\left(bs-ni\right)^{2}\right)\quad\text{as }s\rightarrow ni,$
yielding
$\mathrm{Res}\left[Z_{3},s=\pm\displaystyle\frac{ni}{b}\right]$ $=\displaystyle\frac{\pi^{2}\theta}{b^{2}}\displaystyle\frac{(-1)^{n}\cos(n\theta/b)}{\cosh(an\pi/b)}$ (4.39)
$\quad-\displaystyle\frac{a\pi^{3}}{b^{2}}\displaystyle\frac{(-1)^{n}\sin(n\theta/b)\sinh(an\pi/b)}{\cosh^{2}(an\pi/b)}.$
As $s=0$, the asymptotic behavior near the origin is given by
$Z_{3}(s,\theta;a,b)=\displaystyle\frac{\pi^{2}}{\cos(a\pi s)\cosh(b\pi s)}\left(\displaystyle\frac{\theta}{b^{2}}\displaystyle\frac{1}{s}+\displaystyle\frac{\theta^{3}}{6b^{2}}s-2\zeta(2)\theta s+o(1)\right),\quad s\to 0,$
from which we extract the residue
$\mathrm{Res}\left[Z_{3},s=0\right]=\displaystyle\frac{\theta\pi^{2}}{b^{2}}.$ (4.40)
Applying Lemma 3.1 to the sum of these four residue contributions (4.37)-(4.40) establishes identity (4.27).
The three identities are derived by systematically computing residues of $Z_{1},Z_{2},Z_{3}$ and applying the residue theorem. This completes the proof of Theorem 4.2. ∎
Theorem 4.3.
Let $p>1$ be an odd integer, we have the following equation established
$\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}n^{p}}{\sinh(ny)\cosh^{2}(ny)}$
$\quad\quad\quad\quad\quad\quad\quad=-\displaystyle\frac{\pi^{p+1}}{y^{p+1}}(-1)^{(p-1)/2}\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}n^{p}}{\sinh(\pi^{2}n/y)}$
$\quad\quad\quad\quad\quad\quad\quad\quad+\displaystyle\frac{p\pi^{p}}{2^{p-1}y^{p+1}}(-1)^{(p-1)/2}\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{p-1}}{\sinh((2n-1)\pi^{2}/(2y))}$
$\quad\quad\quad\quad\quad\quad\quad\quad-\displaystyle\frac{\pi^{p+2}}{2^{p}y^{p+2}}(-1)^{(p-1)/2}\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{p}\cosh((2n-1)\pi^{2}/(2y))}{\sinh^{2}((2n-1)\pi^{2}/(2y))},$ (4.41)
$\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{p-1}}{\sinh((2n-1)y/2)\cosh^{2}((2n-1)y/2)}$
$\quad\quad\quad\quad\quad\quad\quad=-\displaystyle\frac{2^{p-1}\pi^{p}}{y^{p}}(-1)^{(p-1)/2}\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}n^{p-1}}{\cosh(\pi^{2}n/y)}$
$\quad\quad\quad\quad\quad\quad\quad\quad+\displaystyle\frac{2(p-1)\pi^{p-1}}{y^{p}}(-1)^{(p-1)/2}\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{p-2}}{\cosh((2n-1)\pi^{2}/(2y))}$
$\quad\quad\quad\quad\quad\quad\quad\quad-\displaystyle\frac{\pi^{p+1}}{y^{p+1}}(-1)^{(p-1)/2}\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{p-1}\sinh((2n-1)\pi^{2}/(2y))}{\cosh^{2}((2n-1)\pi^{2}/(2y))},$ (4.42)
$\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{p}}{\sinh^{2}((2n-1)y/2)\cosh((2n-1)y/2)}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad=-\displaystyle\frac{\pi^{p+1}}{y^{p+1}}(-1)^{(p-1)/2}\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{p}}{\cosh((2n-1)\pi^{2}/(2y))}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-\displaystyle\frac{p2^{p}\pi^{p}}{y^{p+1}}(-1)^{(p-1)/2}\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}n^{p-1}}{\cosh(\pi^{2}n/y)}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+\displaystyle\frac{2^{p}\pi^{p+2}}{y^{p+2}}(-1)^{(p-1)/2}\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}n^{p}\sinh(\pi^{2}n/y)}{\cosh^{2}(\pi^{2}n/y)}.$ (4.43)
Proof.
We prove the first identity in detail, as the proofs for the other two follow similarly. By taking the $p$-th odd derivative of (4.25) in Theorem 4.2 with respect to $\theta$, then $\theta\rightarrow 0$, we obtain the result as follows
$\displaystyle\frac{d^{p}}{d\theta^{p}}\left[\sinh\left(\displaystyle\frac{n\theta}{a}\right)\right]_{|\theta=0}=\left(\displaystyle\frac{n}{a}\right)^{p},\quad\displaystyle\frac{d^{p}}{d\theta^{p}}\left[\theta\cos\left(\displaystyle\frac{2n-1}{2b}\theta\right)\right]_{|\theta=0}=p(-1)^{(p-1)/2}\left(\displaystyle\frac{2n-1}{2b}\right)^{p-1},$
$\displaystyle\frac{d^{p}}{d\theta^{p}}\left[\sin\left(\displaystyle\frac{n\theta}{b}\right)\right]_{|\theta=0}=(-1)^{(p-1)/2}\left(\displaystyle\frac{n}{b}\right)^{p},\quad\displaystyle\frac{d^{p}}{d\theta^{p}}\left[\sin\left(\displaystyle\frac{2n-1}{2b}\theta\right)\right]_{|\theta=0}=(-1)^{(p-1)/2}\left(\displaystyle\frac{2n-1}{2b}\right)^{p}.$
Using the above derivatives and letting $y=\displaystyle\frac{b\pi}{a}$ in (4.25), and we can easily obtain that
$\displaystyle\frac{b}{2}\delta_{p}-(-1)^{(p-1)/2}\displaystyle\frac{ap}{2^{p-1}b^{p-1}}\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{p-1}}{\sinh((2n-1)\pi^{2}/(2y))}$
$\quad\quad\quad+\displaystyle\frac{b^{2}\pi}{a^{p}}\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}n^{p}}{\sinh(ny)\cosh^{2}(ny)}+(-1)^{(p-1)/2}\displaystyle\frac{a\pi}{b^{p-1}}\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}n^{p}}{\sinh(n\pi^{2}/y)}$
$\quad\quad\quad+(-1)^{(p-1)/2}\displaystyle\frac{a^{2}\pi}{2^{p}b^{p}}\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{p}\cosh((2n-1)\pi^{2}/(2y))}{\sinh^{2}((2n-1)\pi^{2}/(2y))}=0,$ (4.44)
where $\delta_{1}=1$ and $\delta_{p}=0$ if $p\geq 3$.
For $p=1$, we obtain the special case
$\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}n}{\sinh(ny)\cosh^{2}(ny)}=$ $-\displaystyle\frac{1}{2y}-\displaystyle\frac{\pi^{2}}{y^{2}}\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}n}{\sinh(\pi^{2}n/y)}$
$+\displaystyle\frac{\pi}{y^{2}}\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}}{\sinh((2n-1)\pi^{2}/(2y))}$
$-\displaystyle\frac{\pi^{3}}{2y^{3}}\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)\cosh((2n-1)\pi^{2}/(2y))}{\sinh^{2}((2n-1)\pi^{2}/(2y))},$ (4.45)
and for $p\geq 3$, we obtain the equation (4.41) .
The proofs of equations (4.42)and (4.43) follow similarly, with the following key differences: For (4.42), we take the $(p-1)$-th order derivative (i.e., even-order derivative) of equation (4.26) with respect to $\theta$; For (4.43), we take the $p$-th order odd derivative of equation (4.27) with respect to $\theta$. ∎
5 Evaluations of Hyperbolic Summations via Jacobi Functions
We now evaluate the following reciprocal hyperbolic series in the Ramanujan type
$G_{p,2}(y)$ $:=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}n^{p}}{\sinh(ny)\cosh^{2}(ny)},$ (5.46)
$G_{p,2}^{\prime}(y)$ $:=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{p-1}}{\sinh((2n-1)y/2)\cosh^{2}((2n-1)y/2)},$ (5.47)
$\overline{G}_{p,2}(y)$ $:=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{p}}{\sinh^{2}((2n-1)y/2)\cosh((2n-1)y/2)}.$ (5.48)
By applying the Fourier series expansions and the Maclaurin series expansions of relevant Jacobi elliptic functions. Let $x,y$ and $z$ satisfy the relations in (2.9).
Theorem 5.1.
Let $p\geq 3$ be an odd integer. We have
$G_{p,2}(y)=$ $(-1)^{(p-1)/2}\displaystyle\frac{(p-1)!z^{p+1}z^{\prime}x(1-x)}{2^{p+1}}\mathrm{R}_{p-1}(1-x)-\displaystyle\frac{pz^{p+1}z^{\prime}x(1-x)}{2^{p}}\mathrm{S}_{p-1}(1-x)\sqrt{1-x}$
$-\displaystyle\frac{z^{p+2}x(1-x)}{2^{2p}}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}\left[\mathrm{S}_{p-1}(1-x)\sqrt{1-x}\right],$ (5.49)
where $\mathrm{R}_{p-1}(1-x)=\displaystyle\frac{(-x)^{(p-3)/2}}{(p-1)!}q_{p-1}\left(\displaystyle\frac{1-x}{-x}\right)\in\mathbb{Q}[x]$ and ${q}_{p-1}\left(x\right)$ represents coefficients in the Maclaurin expansion of $\mathrm{sn}(u)^{2}$; ${\rm S}_{p-1}\left(x\right)$ represents coefficients in the Maclaurin expansion of $\mathrm{cd}\left(u\right)$.
Proof.
Beginning with equation (4.41), we derive
$G_{p,2}(y)$ $=-(-1)^{(p-1)/2}\displaystyle\frac{\pi^{p+1}}{y^{p+1}}X_{p,1}\left(\displaystyle\frac{\pi^{2}}{y}\right)+(-1)^{(p-1)/2}\displaystyle\frac{p\pi^{p}}{2^{p-1}y^{p+1}}Y_{p,1}\left(\displaystyle\frac{\pi^{2}}{y}\right)$
$\quad-(-1)^{(p-1)/2}\displaystyle\frac{\pi^{p+2}}{2^{p}y^{p+2}}\text{DY}_{p,2}\left(\displaystyle\frac{\pi^{2}}{y}\right).$ (5.50)
Now, we need to compute the series $Y_{p,1}\left(y\right)$. According to Lemma 3.5, we obtain the Maclaurin expansion for ${\rm cd}(u)$
$\mathrm{cd}\left(u\right)=\displaystyle\sum_{n=0}^{\infty}{\mathrm{S}_{2n}\left(x\right)\displaystyle\frac{\left(-1\right)^{n}u^{2n}}{\left(2n\right)!}}.$ (5.51)
Using known results from [8], we record the Fourier series expansions of Jacobi elliptic functions ${\rm cd}(u)$
${\rm cd}(u)=\displaystyle\frac{2\pi}{Kk}\displaystyle\sum_{n=0}^{\infty}\displaystyle\frac{(-1)^{n}q^{n+1/2}}{1-q^{2n+1}}\cos\left[(2n+1)\displaystyle\frac{\pi u}{2K}\right]\quad(|q|<1).$ (5.52)
And in the page $165$ of Ramanujan’s Notebooks (III)[5], we get
$\displaystyle\frac{1}{2}z\sqrt{x}{\rm cd}(u)=\displaystyle\sum_{n=0}^{\infty}\displaystyle\frac{(-1)^{n}\cos\left[(2n+1)u\right]}{\sinh((2n+1)y/2)}.$ (5.53)
Applying $q\equiv q\left(x\right)=e^{-y}$, then
${\rm cd}(u)$ $=\displaystyle\frac{2\pi}{\mathrm{K}k}\displaystyle\sum_{n=0}^{\infty}{\displaystyle\frac{(-1)^{n}q^{n+\tfrac{1}{2}}\left(2n+1\right)^{2j}}{1-q^{2n+1}}}\displaystyle\sum_{j=0}^{\infty}{\displaystyle\frac{\left(-1\right)^{j}\left(\displaystyle\frac{\pi u}{2\mathrm{K}}\right)^{2j}}{\left(2j\right)!}}$
$=\displaystyle\frac{\pi}{\mathrm{K}k}\displaystyle\sum_{j=0}^{\infty}{\displaystyle\frac{\left(-1\right)^{j}\left(\displaystyle\frac{\pi u}{2\mathrm{K}}\right)^{2j}}{\left(2j\right)!}}Y_{2j+1,1}\left(y\right).$ (5.54)
Comparing the coefficients of $u^{2n}$ in (5.51) and (5), we deduce
$Y_{2n+1,1}\left(y\right)=\displaystyle\sum_{j=1}^{\infty}{\displaystyle\frac{(-1)^{j}\left(2j-1\right)^{2n}}{\sinh\left(\displaystyle\frac{2j-1}{2}y\right)}}=(-1)^{n}z^{2n+1}\sqrt{x}\displaystyle\frac{\mathrm{S}_{2n+1}\left(x\right)}{2}.$ (5.55)
Differentiating (5.55) gives the relation
$\displaystyle\frac{\mathrm{d}}{\mathrm{d}y}Y_{p,1}\left(y\right)=\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{\mathrm{d}}{\mathrm{d}y}\displaystyle\frac{(-1)^{n}\left(2n-1\right)^{p-1}}{\sinh\left(\displaystyle\frac{2n-1}{2}y\right)}=-\displaystyle\frac{1}{2}}\text{DY}_{p,2}\left(y\right).$
Using the transformation $\frac{dx}{dy}=-x(1-x)z^{2}$ ([5, P. 120, Entry. 9(i)] or (2.11)), we obtain
$\text{DY}_{p,2}\left(y\right)=2x\left(1-x\right)z^{2}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}Y_{p,1}\left(y\right).$ (5.56)
Applying Lemma 3.2, for $\Omega\left(x,e^{-y},z,z^{\prime}\right)=0$, we have
$\Omega\left(1-x,e^{-\pi^{2}/y},yz/\pi,\displaystyle\frac{1}{\pi}\left(\displaystyle\frac{1}{x\left(1-x\right)z}-yz^{\prime}\right)\right)=0.$
Hence, the (5.55) and (5.56) can be rewritten as
$Y_{p,1}\left(\displaystyle\frac{\pi^{2}}{y}\right)=-(-1)^{(p-1)/2}\left(\displaystyle\frac{yz}{\pi}\right)^{p}\displaystyle\frac{\mathrm{S}_{p-1}\left(1-x\right)}{2}\sqrt{1-x},$ (5.57)
and
$\text{DY}_{p,2}\left(\displaystyle\frac{\pi^{2}}{y}\right)$ $=2x\left(1-x\right)\left(\displaystyle\frac{yz}{\pi}\right)^{2}\displaystyle\frac{\mathrm{d}}{\mathrm{d}\left(1-x\right)}Y_{p,1}\left(\displaystyle\frac{\pi^{2}}{y}\right)$
$=-2x\left(1-x\right)\left(\displaystyle\frac{yz}{\pi}\right)^{2}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}Y_{p,1}^{\prime}\left(\displaystyle\frac{\pi^{2}}{y}\right).$ (5.58)
What’s more, as known from Rui-Xu-Zhao’s paper[17, Thm. 3.15] and apply Lemma 3.2, for $\Omega\left(x,e^{-y},z,z^{\prime}\right)=0$, we have
$\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}n^{p}}{\sinh\left(\displaystyle\frac{\pi^{2}}{y}\right)}=-\displaystyle\frac{(p-1)!}{2^{p+1}}\left(\displaystyle\frac{yz}{\pi}\right)^{2}x(1-x)\mathrm{R}_{p-1}(1-x),$ (5.59)
where $\mathrm{R}_{p-1}(1-x)=\displaystyle\frac{(-x)^{(p-3)/2}}{(p-1)!}q_{p-1}\left(\displaystyle\frac{1-x}{-x}\right)\in\mathbb{Q}[x]$.
Finally, combining equations (5), (5.55), (5.57), (5) and (5.59), we thus complete the proof of Theorem 5.1. ∎
Example 5.2.
By Mathematica, we have
$\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}n^{3}}{\sinh(ny)\cosh^{2}(ny)}=\displaystyle\frac{1}{16}xz^{4}\left\{6z^{\prime}x(1-x)^{3/2}-\sqrt{1-x}(3x-2)z+2(x-1)\right\},$
$\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}n^{7}}{\sinh(ny)\cosh^{2}(ny)}$
$=\displaystyle\frac{1}{256}xz^{8}\Big{\{}14z^{\prime}x\left(61x^{2}-76x+16\right)(1-x)^{3/2}\Big{.}+16(x-1)\left(17x^{2}-17x+2\right)\atop-\sqrt{1-x}(427x^{3}-746x^{2}+352x-32)z\Big{\}},$
$\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}n^{11}}{\sinh(ny)\cosh^{2}(ny)}$
$=\displaystyle\frac{1}{4096}xz^{12}\Big{\{}22z^{\prime}x\left(50521x^{4}-113672x^{3}+79728x^{2}-16832x+256\right)(1-x)^{3/2}\Big{.}$
.$\Big{.}\qquad\qquad\qquad+256(x-1)\left(1382x^{4}-2764x^{3}+1641x^{2}-259x+2\right)\Big{.}$
.$\Big{.}\quad-\sqrt{1-x}\left(555731x^{5}-1528258x^{4}+1467472x^{3}-562528x^{2}+68096x-512\right)z\Big{\}},$
$\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}n^{15}}{\sinh(ny)\cosh^{2}(ny)}$
$=\displaystyle\frac{1}{65536}xz^{16}\Big{\{}30z^{\prime}x\left(199360981x^{6}-647923188x^{5}+775638816x^{4}\atop-408850432x^{3}+85975296x^{2}-4205568x+4096\right)(1-x)^{3/2}\Big{.}$
.$\Big{.}\quad\qquad\qquad\qquad+2048(x-1)\left(929569x^{6}-2788707x^{5}+3021099x^{4}\atop-1394353x^{3}+240594x^{2}-8202x+4\right)\Big{.}$
.$\Big{.}\quad-\sqrt{1-x}\left(2990414715x^{7}-11214055178x^{6}+16307105232x^{5}-11436042048x^{4}\atop+3872630528x^{3}-536879616x^{2}+16834560x-8192\right)z\Big{\}}.$
Corollary 5.3.
Set $\Gamma=\Gamma(1/4)$ and $p=4m-1$. Then, for any integer $m>0$, we know that
$\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}n^{4m-1}}{\sinh(n\pi)\cosh^{2}(n\pi)}$ $=-\displaystyle\frac{1}{2^{8m+2}}(4m-2)!\mathrm{R}_{4m-2}\left(\displaystyle\frac{1}{2}\right)\displaystyle\frac{\Gamma^{8m}}{\pi^{6m}}$
$\quad-\displaystyle\frac{1}{2^{8m-1}}(4m-1)\mathrm{S}_{4m-2}\left(\displaystyle\frac{1}{2}\right)\displaystyle\frac{1}{\sqrt{2}}\displaystyle\frac{\Gamma^{8m-2}}{\pi^{(12m-1)/2}}$
$\quad-\displaystyle\frac{1}{2^{8m+2}}\left(\mathrm{S}_{4m-2}^{\prime}\left(\displaystyle\frac{1}{2}\right)-\mathrm{S}_{4m-2}\left(\displaystyle\frac{1}{2}\right)\right)\displaystyle\frac{1}{\sqrt{2}}\displaystyle\frac{\Gamma^{8m+2}}{\pi^{(12m+3)/2}}.$ (5.60)
Example 5.4.
Set $\Gamma=\Gamma(1/4)$. By Mathematica, we have
$\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}n^{3}}{\sinh(n\pi)\cosh^{2}(n\pi)}$ $=\displaystyle\frac{\Gamma^{10}}{2048\sqrt{2}\pi^{15/2}}-\displaystyle\frac{\Gamma^{8}}{512\pi^{6}}+\displaystyle\frac{3\Gamma^{6}}{256\sqrt{2}\pi^{11/2}},$
$\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}n^{7}}{\sinh(n\pi)\cosh^{2}(n\pi)}$ $=-\displaystyle\frac{87\Gamma^{18}}{2097152\sqrt{2}\pi^{27/2}}+\displaystyle\frac{9\Gamma^{16}}{65536\pi^{12}}-\displaystyle\frac{189\Gamma^{14}}{262144\sqrt{2}\pi^{23/2}},$
$\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}n^{11}}{\sinh(n\pi)\cosh^{2}(n\pi)}$ $=\displaystyle\frac{57969\Gamma^{26}}{2147483648\sqrt{2}\pi^{39/2}}-\displaystyle\frac{189\Gamma^{24}}{2097152\pi^{18}}+\displaystyle\frac{126819\Gamma^{22}}{268435456\sqrt{2}\pi^{35/2}},$
$\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}n^{15}}{\sinh(n\pi)\cosh^{2}(n\pi)}$ $=-\displaystyle\frac{160692903\Gamma^{34}}{2199023255552\sqrt{2}\pi^{51/2}}+\displaystyle\frac{130977\Gamma^{32}}{536870912\pi^{24}}-\displaystyle\frac{351673245\Gamma^{30}}{274877906944\sqrt{2}\pi^{47/2}}.$
Theorem 5.5.
Let $p\geq 3$ be an odd integer. We have
$G_{p,2}^{\prime}\left(y\right)$ $=-\displaystyle\frac{1}{2}z^{p}\mathrm{A}_{p-1}(1-x)\sqrt{x}+(p-1)z^{p}z^{\prime}x(1-x)\mathrm{P}_{p-2}(1-x)\sqrt{x(1-x)}$
$\quad+x(1-x)z^{p+1}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}\left[\mathrm{P}_{p-2}(1-x)\sqrt{x(1-x)}\right],$ (5.61)
where $\mathrm{A}_{p-1}\left(1-x\right)$ are coefficients in the Maclaurin expansion of $\mathrm{nd}\left(u\right)$ and $\mathrm{P}_{p-2}\left(1-x\right)$
are coefficients in the Maclaurin expansion of $\mathrm{sd}\left(u\right)$.
Proof.
Starting from equation (4.42), we derive
$G_{p,2}^{\prime}\left(y\right)$ $=-(-1)^{(p-1)/2}\displaystyle\frac{2^{p-1}\pi^{p}}{y^{p}}B_{p,1}\left(\displaystyle\frac{\pi^{2}}{y}\right)+(-1)^{(p-1)/2}\displaystyle\frac{2(p-1)\pi^{p-1}}{y^{p}}X_{p-2,1}^{\prime}\left(\displaystyle\frac{\pi^{2}}{y}\right)$
$\quad-(-1)^{(p-1)/2}\displaystyle\frac{\pi^{p+1}}{y^{p+1}}\text{DX}_{p-2,2}^{\prime}\left(\displaystyle\frac{\pi^{2}}{y}\right).$ (5.62)
Now, we need to compute the series $B_{p,1}\left(y\right)$. From Lemma 3.5 and [8], we obtain the Maclaurin expansion for $nd(u)$
$\mathrm{nd}(u)=\displaystyle\sum_{n=0}^{\infty}{\mathrm{A}_{2n}\left(x\right)\displaystyle\frac{\left(-1\right)^{n}u^{2n}}{\left(2n\right)!}},\quad\displaystyle\frac{2q^{n}}{1+q^{2n}}=\displaystyle\frac{1}{\cosh(ny)}.$ (5.63)
and its Fourier series representation
${\rm nd}(u)=\displaystyle\frac{\pi}{2Kk^{\prime}}\displaystyle\frac{2\pi}{Kk^{\prime}}\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}q^{n}}{1-q^{2n+1}}\cos\left[(2n)\displaystyle\frac{\pi u}{2K}\right]\quad(|q|<1).$ (5.64)
Setting $q\equiv q\left(x\right)=e^{-y}$, we expand
${\rm nd}(u)$ $=\displaystyle\frac{\pi}{2Kk^{\prime}}+\displaystyle\frac{2\pi}{\mathrm{K}k^{\prime}}\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{(-1)^{n}q^{n}n^{2j}}{1+q^{2n}}}\displaystyle\sum_{j=0}^{\infty}{\displaystyle\frac{\left(-1\right)^{j}\left(\displaystyle\frac{\pi u}{\mathrm{K}}\right)^{2j}}{\left(2j\right)!}}$
$=\displaystyle\frac{\pi}{2\mathrm{K}k^{\prime}}+\displaystyle\frac{\pi}{\mathrm{K}k^{\prime}}\displaystyle\sum_{j=0}^{\infty}{\displaystyle\frac{\left(-1\right)^{j}\left(\displaystyle\frac{\pi u}{\mathrm{K}}\right)^{2j}}{\left(2j\right)!}}B_{2j+1,1}\left(y\right).$ (5.65)
Comparing the coefficients of $u^{2n}$ in (5.63) and (5), we deduce
$B_{p,1}\left(y\right)=(-1)^{(p-1)/2}z^{p}\sqrt{1-x}\displaystyle\frac{\mathrm{A}_{p-1}\left(x\right)}{2^{p}}.$ (5.66)
Applying Lemma 3.2, for $\Omega\left(x,e^{-y},z,z^{\prime}\right)=0$, we obtain
$B_{p,1}\left(\displaystyle\frac{\pi^{2}}{y}\right)=(-1)^{(p-1)/2}\left(\displaystyle\frac{yz}{\pi}\right)^{p}\sqrt{x}\displaystyle\frac{\mathrm{A}_{p-1}\left(1-x\right)}{2^{p}}.$ (5.67)
By using the results in Rui-Xu-Zhao’s paper[17, Thm. 3.3], we have
$X_{p,1}^{\prime}(y)=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{p}}{\cosh\left(\displaystyle\frac{(2n-1)y}{2}\right)}=-(-1)^{(p-1)/2}z^{p+1}\sqrt{x(1-x)}\displaystyle\frac{\mathrm{P}_{p}(x)}{2}.$ (5.68)
Differentiating $X_{p-2,1}^{\prime}\left(y\right)$ yields
$\displaystyle\frac{\mathrm{d}}{\mathrm{d}y}X_{p-2,1}^{\prime}\left(y\right)=\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{\mathrm{d}}{\mathrm{d}y}\displaystyle\frac{(-1)^{n}\left(2n-1\right)^{p-2}}{\cosh\left(\displaystyle\frac{2n-1}{2}y\right)}=-\displaystyle\frac{1}{2}}\text{DX}_{p-2,2}^{\prime}\left(y\right).$
Using $\frac{dx}{dy}=-x(1-x)z^{2}$ ([5, P. 120, Entry. 9(i)] or (2.11)), we get
$\text{DX}_{p-2,2}^{\prime}\left(y\right)=2x\left(1-x\right)z^{2}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}X_{p-2,1}^{\prime}\left(y\right).$ (5.69)
Applying Lemma 3.2, for $\Omega\left(x,e^{-y},z,z^{\prime}\right)=0$, we obtain
$\Omega\left(1-x,e^{-\pi^{2}/y},yz/\pi,\displaystyle\frac{1}{\pi}\left(\displaystyle\frac{1}{x\left(1-x\right)z}-yz^{\prime}\right)\right)=0.$
Hence, the (5.68) and (5.69) can be rewritten as
$X_{p-2,1}^{\prime}\left(\displaystyle\frac{\pi^{2}}{y}\right)=(-1)^{(p-1)/2}\left(\displaystyle\frac{yz}{\pi}\right)^{p+1}\displaystyle\frac{\mathrm{P}_{p-2}\left(1-x\right)}{2}\sqrt{x(1-x)}$ (5.70)
and
$\text{DX}_{p-2,2}^{\prime}\left(\displaystyle\frac{\pi^{2}}{y}\right)$ $=2x\left(1-x\right)\left(\displaystyle\frac{yz}{\pi}\right)^{2}\displaystyle\frac{\mathrm{d}}{\mathrm{d}\left(1-x\right)}X_{p-2,1}^{\prime}\left(\displaystyle\frac{\pi^{2}}{y}\right)$
$=-2x\left(1-x\right)\left(\displaystyle\frac{yz}{\pi}\right)^{2}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}X_{p-2,1}^{\prime}\left(\displaystyle\frac{\pi^{2}}{y}\right).$ (5.71)
Finally, substituting (5.67), (5.70) and (5) into (5) completes the proof through direct calculation. ∎
Example 5.6.
By Mathematica, we have
$\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{2}}{\sinh((2n-1)y/2)\cosh^{2}((2n-1)y/2)}$
$=\displaystyle\frac{1}{2}z^{3}\sqrt{(1-x)x}\Big{\{}4z^{\prime}(1-x)x+(1-2x)z\Big{.}-\sqrt{(1-x)}\Big{\}},$
$\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{6}}{\sinh((2n-1)y/2)\cosh^{2}((2n-1)y/2)}$
$=\displaystyle\frac{1}{2}z^{7}\sqrt{x}\Big{\{}12z^{\prime}\left(16x^{2}-16x+1\right)(1-x)^{3/2}x\Big{.}+\left(-96x^{3}+144x^{2}-50x+1\right)\sqrt{1-x}z\Big{.}$
.$\Big{.}\quad+61x^{3}-107x^{2}+47x-1\Big{\}},$
$\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{10}}{\sinh((2n-1)y/2)\cosh^{2}((2n-1)y/2)}$
$=\displaystyle\frac{1}{2}z^{11}\sqrt{x}\Big{\{}20z^{\prime}\left(7936x^{4}-15872x^{3}+9168x^{2}-1232x+1\right)(1-x)^{3/2}x\Big{.}$
.$\Big{.}\quad+\left(-79360x^{5}+198400x^{4}-166112x^{3}+50768x^{2}-3698x+1\right)\sqrt{1-x}z\Big{.}$
.$\Big{.}\quad+50521x^{5}-138933x^{4}+130250x^{3}-45530x^{2}+3693x-1\Big{\}},$
$\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{14}}{\sinh((2n-1)y/2)\cosh^{2}((2n-1)y/2)}$
$=\displaystyle\frac{1}{2}z^{15}\sqrt{x}\Big{\{}28z^{\prime}\left(22368256x^{6}-67104768x^{5}+71997696x^{4}\atop-32154112x^{3}+4992576x^{2}-99648x+1\right)(1-x)^{3/2}x\Big{.}$
.$\Big{.}\quad+\left(-313155584x^{7}+1096044544x^{6}-1458129408x^{5}\atop+905212160x^{4}-255034240x^{3}+25361472x^{2}-298946x+1\right)\sqrt{(1-x)}z\Big{.}$
.$\Big{.}\quad+(199360981x^{7}-747603679x^{6}+1074680289x^{5}-728130163x^{4}\atop+226132303x^{3}-24738669x^{2}+298939x-1\Big{\}}.$
Corollary 5.7.
Set $\Gamma=\Gamma(1/4)$ and $p=4m-1$. Then, for any integer $m>0$ we know that
$\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{4m-2}}{\sinh((2n-1)\pi/2)\cosh^{2}((2n-1)\pi/2)}$ $=-\displaystyle\frac{1}{2^{4m}}\mathrm{A}_{4m-2}\left(\displaystyle\frac{1}{2}\right)\displaystyle\frac{1}{\sqrt{2}}\displaystyle\frac{\Gamma^{8m-2}}{\pi^{(12m-3)/2}}$
$\quad-\displaystyle\frac{1}{2^{4m}}(4m-2)\mathrm{P}_{4m-3}\left(\displaystyle\frac{1}{2}\right)\displaystyle\frac{1}{\sqrt{2}}\displaystyle\frac{\Gamma^{8m-4}}{\pi^{6m-2}}.$ (5.72)
Example 5.8.
Set $\Gamma=\Gamma(1/4)$. By Mathematica, we have
$\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{2}}{\sinh((2n-1)\pi/2)\cosh^{2}((2n-1)\pi/2)}$ $=\displaystyle\frac{\Gamma^{4}}{8\pi^{4}}-\displaystyle\frac{\Gamma^{6}}{32\sqrt{2}\pi^{9/2}},$
$\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{6}}{\sinh((2n-1)\pi/2)\cosh^{2}((2n-1)\pi/2)}$ $=\displaystyle\frac{27\Gamma^{14}}{2048\sqrt{2}\pi^{21/2}}-\displaystyle\frac{9\Gamma^{12}}{128\pi^{10}},$
$\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{10}}{\sinh((2n-1)\pi/2)\cosh^{2}((2n-1)\pi/2)}$ $=\displaystyle\frac{945\Gamma^{20}}{2048\pi^{16}}-\displaystyle\frac{11529\Gamma^{22}}{131072\sqrt{2}\pi^{33/2}},$
$\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{14}}{\sinh((2n-1)\pi/2)\cosh^{2}((2n-1)\pi/2)}$ $=\displaystyle\frac{23444883\Gamma^{30}}{8388608\sqrt{2}\pi^{45/2}}-\displaystyle\frac{480249\Gamma^{28}}{32768\pi^{22}}.$
Theorem 5.9.
Let $p\geq 3$ be an odd integer. We have
$\bar{G}_{p,1}\left(y\right)$ $=\displaystyle\frac{1}{2}z^{p+1}\mathrm{P}_{p-1}(1-x)\sqrt{x(1-x)}+pz^{p+1}z^{\prime}x(1-x)\mathrm{A}_{p-1}(1-x)\sqrt{x}$ (5.73)
$\quad-x(1-x)z^{p+2}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}\left[\mathrm{A}_{p-1}(1-x)\sqrt{x}\right],$
where $\mathrm{P}_{p}\left(1-x\right)$ denotes coefficients in the Maclaurin expansion of $\mathrm{sd}\left(u\right)$ and $\mathrm{A}_{p-1}\left(1-x\right)$
denotes coefficients in the Maclaurin expansion of $\mathrm{nd}\left(u\right)$.
Proof.
We begin with equation (4.43), which gives
$\bar{G}_{p,1}\left(y\right)$ $=-(-1)^{(p-1)/2}\displaystyle\frac{\pi^{p+1}}{y^{p+1}}X_{p,1}^{\prime}\left(\displaystyle\frac{\pi^{2}}{y}\right)-(-1)^{(p-1)/2}\displaystyle\frac{p2^{p}\pi^{p}}{y^{p}}B_{p,1}\left(\displaystyle\frac{\pi^{2}}{y}\right)$
$\quad+(-1)^{(p-1)/2}\displaystyle\frac{2^{p}\pi^{p+2}}{y^{p+2}}\text{DB}_{p,2}\left(\displaystyle\frac{\pi^{2}}{y}\right).$ (5.74)
The derivative of $B_{p,1}^{\prime}\left(y\right)$ satisfies
$\displaystyle\frac{\mathrm{d}}{\mathrm{d}y}B_{p,1}\left(y\right)=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{\mathrm{d}}{\mathrm{d}y}\displaystyle\frac{(-1)^{n}n^{p-1}}{\cosh(ny)}=-\text{DB}_{p,2}\left(y\right).$
From the fundamental relation $\frac{dx}{dy}=-x(1-x)z^{2}$ ([5, P. 120, Entry. 9(i)] or [17]), we have
$\text{DB}_{p,2}\left(y\right)=x\left(1-x\right)z^{2}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}B_{p,1}\left(y\right).$ (5.75)
Using Lemma 3.2, for $\Omega\left(x,e^{-y},z,z^{\prime}\right)=0$, we obtain
$\Omega\left(1-x,e^{-\pi^{2}/y},yz/\pi,\displaystyle\frac{1}{\pi}\left(\displaystyle\frac{1}{x\left(1-x\right)z}-yz^{\prime}\right)\right)=0.$
In addition, it is obvious to obtain from (5.67) and (5.70) that
$X_{p,1}^{\prime}\left(\displaystyle\frac{\pi^{2}}{y}\right)=-(-1)^{(p-1)/2}\left(\displaystyle\frac{yz}{\pi}\right)^{p+1}\displaystyle\frac{\mathrm{P}_{p}\left(1-x\right)}{2}\sqrt{x(1-x)}$ (5.76)
and
$\text{DB}_{p,2}\left(\displaystyle\frac{\pi^{2}}{y}\right)$ $=x\left(1-x\right)\left(\displaystyle\frac{yz}{\pi}\right)^{2}\displaystyle\frac{\mathrm{d}}{\mathrm{d}\left(1-x\right)}B_{p,1}\left(\displaystyle\frac{\pi^{2}}{y}\right)$
$=-x\left(1-x\right)\left(\displaystyle\frac{yz}{\pi}\right)^{2}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}B_{p,1}\left(\displaystyle\frac{\pi^{2}}{y}\right).$ (5.77)
Finally, substituting (5.67), (5.76) and (5) into (5) and performing the necessary calculations completes the proof. ∎
Example 5.10.
By Mathematica, we have
$\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{(-1)^{n}(2n-1)^{3}}{\sinh^{2}((2n-1)y/2)\cosh\mathrm{((}2n-1)y/2)}}$
$=\displaystyle\frac{1}{2}\sqrt{x}z^{4}\{-6z^{\prime}(x-1)^{2}x-(x-1)(3x-1)z.+(1-2x)\sqrt{1-x}\},$
$\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{(-1)^{n}(2n-1)^{7}}{\sinh^{2}((2n-1)y/2)\cosh\mathrm{((}2n-1)y/2)}}$
$=\displaystyle\frac{1}{2}\sqrt{x}z^{8}\left\{\begin{array}[]{c}\left(-272x^{3}+408x^{2}-138x+1\right)\sqrt{1-x}\\ -14z^{\prime}(x-1)\left(61x^{3}-107x^{2}+47x-1\right)x\\ -(x-1)\left(427x^{3}-535x^{2}+141x-1\right)z\\ \end{array}\right\},$
$\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{(-1)^{n}(2n-1)^{11}}{\sinh^{2}((2n-1)y/2)\cosh\mathrm{((}2n-1)y/2)}}$
$=\displaystyle\frac{1}{2}\sqrt{x}z^{12}\left\{\begin{array}[]{c}\left(-353792x^{5}+884480x^{4}-729728x^{3}+210112x^{2}-11074x+1\right)\sqrt{1-x}\\ -(x-1)\left(555731x^{5}-1250397x^{4}+911750x^{3}-227650x^{2}+11079x-1\right)z\\ -22z^{\prime}(x-1)\left(50521x^{5}-138933x^{4}+130250x^{3}-45530x^{2}+3693x-1\right)x\\ \end{array}\right\},$
$\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{(-1)^{n}(2n-1)^{15}}{\sinh^{2}((2n-1)y/2)\cosh\mathrm{((}2n-1)y/2)}}$
$=\displaystyle\frac{1}{2}\sqrt{x}z^{16}\left\{\begin{array}[]{c}\left(\begin{array}[]{c}-1903757312x^{7}+6663150592x^{6}-8804878080x^{5}+5354318720x^{4}\\ -1429612624x^{3}+121675512x^{2}-896810x+1\\ \end{array}\right)\sqrt{1-x}\\ -(x-1)\left(\begin{array}[]{c}2990414715x^{7}-9718847827x^{6}+11821483179x^{5}\\ -6553171467x^{4}+1582926121x^{3}-123693345x^{2}+896817x-1\\ \end{array}\right)z\\ -30z^{\prime}(x-1)\left(\begin{array}[]{c}199360981x^{7}-747603679x^{6}+1074680289x^{5}-728130163x^{4}\\ +226132303x^{3}-24738669x^{2}+298939x-1\\ \end{array}\right)x\\ \end{array}\right\}.$
Corollary 5.11.
Set $\Gamma=\Gamma(1/4)$ and $p=4m-1$. Then, for any integer $m>0$, we know that
$\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{4m-1}}{\sinh^{2}((2n-1)\pi/2)\cosh((2n-1)\pi/2)}$
$=-\displaystyle\frac{1}{2^{4m}}(4m-1)\mathrm{A}_{4m-2}\left(\displaystyle\frac{1}{2}\right)\displaystyle\frac{1}{\sqrt{2}}\displaystyle\frac{\Gamma^{8m-2}}{\pi^{(12m-1)/2}}$
$\quad-\displaystyle\frac{1}{2^{4m+3}}(4m-2)\left(\mathrm{A}_{4m-2}\left(\displaystyle\frac{1}{2}\right)^{\prime}+\mathrm{A}_{4m-2}\left(\displaystyle\frac{1}{2}\right)\right)\displaystyle\frac{1}{\sqrt{2}}\displaystyle\frac{\Gamma^{8m+2}}{\pi^{(12m+3)/2}}.$ (5.78)
Example 5.12.
Set $\Gamma=\Gamma(1/4)$. By Mathematica, we have
$\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{3}}{\sinh^{2}((2n-1)\pi/2)\cosh((2n-1)\pi/2)}$ $=\displaystyle\frac{\Gamma^{10}}{256\sqrt{2}\pi^{15/2}}-\displaystyle\frac{3\Gamma^{6}}{32\sqrt{2}\pi^{11/2}},$
$\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{7}}{\sinh^{2}((2n-1)\pi/2)\cosh((2n-1)\pi/2)}$ $=\displaystyle\frac{189\Gamma^{14}}{2048\sqrt{2}\pi^{23/2}}-\displaystyle\frac{87\Gamma^{18}}{16384\sqrt{2}\pi^{27/2}},$
$\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{11}}{\sinh^{2}((2n-1)\pi/2)\cosh((2n-1)\pi/2)}$ $=\displaystyle\frac{57969\Gamma^{26}}{1048576\sqrt{2}\pi^{39/2}}-\displaystyle\frac{126819\Gamma^{22}}{131072\sqrt{2}\pi^{35/2}},$
$\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{15}}{\sinh^{2}((2n-1)\pi/2)\cosh((2n-1)\pi/2)}$ $=\displaystyle\frac{351673245\Gamma^{30}}{8388608\sqrt{2}\pi^{47/2}}-\displaystyle\frac{160692903\Gamma^{34}}{67108864\sqrt{2}\pi^{51/2}}.$
6 Evaluations of Berndt-type Integrals
Theorem 6.1.
Let $m\in\mathbb{N}$ and $\Gamma=\Gamma(1/4)$. Then, the following integral evaluates as
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m-1}\mathrm{d}x}{\left[\cosh(2x)-\cos(2x)\right]\left[\cosh x+\cos x\right]}}$ $=c_{1,m}\displaystyle\frac{\Gamma^{8m-4}}{\pi^{2m-1}}+\displaystyle\frac{c_{2,m}}{\sqrt{2}}\displaystyle\frac{\Gamma^{8m-2}}{\pi^{(4m-1)/2}}$ (6.79)
$\quad+c_{3,m}\displaystyle\frac{\Gamma^{8m}}{\pi^{2m}}+\displaystyle\frac{c_{4,m}}{\sqrt{2}}\displaystyle\frac{\Gamma^{8m+2}}{\pi^{(4m+3)/2}}+c_{5,m}\displaystyle\frac{\Gamma^{8m+4}}{\pi^{2m+3}},$
where the constants $c_{1,m},c_{2,m},c_{3,m},c_{4,m},c_{5,m}\in\mathbb{Q}$.
Proof.
Following the work of Rui-Xu-Zhao’s[17], let $\Gamma=\Gamma(1/4)$, and for positive integers $n$ and $k$, define $\mathrm{P}_{n}^{(k)}:=\mathrm{P}_{n}^{(k)}(1/2)$, $\mathrm{S}_{n}^{(k)}:=\mathrm{S}_{n}^{(k)}(1/2)$ and $\mathrm{A}_{n}^{(k)}=\mathrm{A}_{n}^{(k)}(1/2)$.
Besides, $\mathrm{P}_{4m-1}(1/2)=\mathrm{P}_{4m-3}^{\prime}(1/2)=0$. Then, for positive integer $m>0$ we know that
$\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{4m-1}}{\cosh^{3}\left(\displaystyle\frac{(2n-1)\pi}{2}\right)}=\displaystyle\frac{\Gamma^{8m-4}}{2^{4m+7}\pi^{6m+3}}\left\{\begin{array}[]{l}128(8m^{2}-6m+1)\pi^{4}p_{4m-3}\\ +\Gamma^{8}[(4m-6)p_{4m-3}+p_{4m-3}^{\prime\prime}]\end{array}\right\}.$ (6.82)
By substituting the equations (5.3), (5.7), (5.11) and (6.82) into Theorem 4.1, we obtain
$2\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m-1}\mathrm{d}x}{\left[\cosh(2x)-\cos(2x)\right]\left[\cosh x+\cos x\right]}}$
$=(-1)^{m-1}\displaystyle\frac{1}{2^{6m+1}}(8m^{2}-6m+1)\mathrm{P}_{4m-3}\displaystyle\frac{\Gamma^{8m-4}}{\pi^{2m-1}}$
$\quad+(-1)^{m-1}\displaystyle\frac{1}{2^{6m+2}}(4m-1)(\mathrm{S}_{4m-2}-\mathrm{A}_{4m-2})\displaystyle\frac{1}{\sqrt{2}}\displaystyle\frac{\Gamma^{8m-2}}{\pi^{(4m-1)/2}}$
$\quad+(-1)^{m-1}\displaystyle\frac{1}{2^{8m+3}}q_{4m-2}(-1)\displaystyle\frac{\Gamma^{8m}}{\pi^{2m}}$
$\quad+(-1)^{m-1}\displaystyle\frac{1}{2^{6m+5}}(\mathrm{S}_{4m-2}^{\prime}-\mathrm{S}_{4m-2}+\mathrm{A}_{4m-2}^{\prime}+\mathrm{A}_{4m-2})\displaystyle\frac{1}{\sqrt{2}}\displaystyle\frac{\Gamma^{8m-2}}{\pi^{(4m+3)/2}}$
$\quad+(-1)^{m}\displaystyle\frac{1}{2^{6m+8}}\left[(4m-6)\mathrm{P}_{4m-3}+\mathrm{P}_{4m-3}^{\prime\prime}\right]\displaystyle\frac{\Gamma^{8m+4}}{\pi^{2m+3}}.$ (6.83)
From this, the coefficients $c_{i,m}$ can be explicitly identified as
$c_{1,m}$ $=(-1)^{m-1}\displaystyle\frac{1}{2^{6m+2}}(8m^{2}-6m+1)\mathrm{P}_{4m-3}\in\mathbb{Q},$ (6.84)
$c_{2,m}$ $=(-1)^{m-1}\displaystyle\frac{1}{2^{6m+3}}(4m-1)(\mathrm{S}_{4m-2}-\mathrm{A}_{4m-2})\in\mathbb{Q},$ (6.85)
$c_{3,m}$ $=(-1)^{m-1}\displaystyle\frac{1}{2^{8m+4}}q_{4m-2}(-1)\in\mathbb{Q},$ (6.86)
$c_{4,m}$ $=(-1)^{m-1}\displaystyle\frac{1}{2^{6m+6}}(\mathrm{S}_{4m-2}^{\prime}-\mathrm{S}_{4m-2}+\mathrm{A}_{4m-2}^{\prime}+\mathrm{A}_{4m-2})\in\mathbb{Q},$ (6.87)
$c_{5,m}$ $=(-1)^{m}\displaystyle\frac{1}{2^{6m+9}}\left[(4m-6)\mathrm{P}_{4m-3}+\mathrm{P}_{4m-3}^{\prime\prime}\right]\in\mathbb{Q}.$ (6.88)
Thus, we have completed the proof of this theorem. ∎
With Mathematica, we can compute the following using the formulas from Theorem 6.1.
Example 6.2.
Let $\Gamma=\Gamma(1/4)$. We have
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{3}\mathrm{d}x}{\left[\cosh(2x)-\cos(2x)\right]\left[\cosh x+\cos x\right]}}$ $=\displaystyle\frac{\Gamma^{12}}{16384\pi^{5}}-\displaystyle\frac{\Gamma^{10}}{4096\sqrt{2}\pi^{7/2}}+\displaystyle\frac{\Gamma^{8}}{2048\pi^{2}}$
$\quad-\displaystyle\frac{3\Gamma^{6}}{512\sqrt{2}\pi^{3/2}}+\displaystyle\frac{3\Gamma^{4}}{256\pi},$
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{7}\mathrm{d}x}{\left[\cosh(2x)-\cos(2x)\right]\left[\cosh x+\cos x\right]}}$ $=\displaystyle\frac{13\Gamma^{20}}{1048576\pi^{7}}-\displaystyle\frac{87\Gamma^{18}}{1048576\sqrt{2}\pi^{11/2}}+\displaystyle\frac{9\Gamma^{16}}{65536\pi^{4}}$
$\quad-\displaystyle\frac{189\Gamma^{14}}{131072\sqrt{2}\pi^{7/2}}+\displaystyle\frac{63\Gamma^{11}}{16384\pi^{3}},$
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{11}\mathrm{d}x}{\left[\cosh(2x)-\cos(2x)\right]\left[\cosh x+\cos x\right]}}$ $=\displaystyle\frac{2169\Gamma^{28}}{67108864\pi^{9}}-\displaystyle\frac{57969\Gamma^{26}}{268435456\sqrt{2}\pi^{15/2}}+\displaystyle\frac{189\Gamma^{24}}{524288\pi^{6}}$
$\quad-\displaystyle\frac{126819\Gamma^{22}}{33554432\sqrt{2}\pi^{11/2}}+\displaystyle\frac{10395\Gamma^{20}}{1048576\pi^{5}},$
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{15}\mathrm{d}x}{\left[\cosh(2x)-\cos(2x)\right]\left[\cosh x+\cos x\right]}}$ $=\displaystyle\frac{1504197\Gamma^{36}}{4294967296\pi^{11}}-\displaystyle\frac{160692903\Gamma^{34}}{68719476736\sqrt{2}\pi^{19/2}}$
$\quad+\displaystyle\frac{130977\Gamma^{32}}{33554432\pi^{8}}-\displaystyle\frac{351673245\Gamma^{30}}{8589934592\sqrt{2}\pi^{15/2}}$
$\quad+\displaystyle\frac{7203735\Gamma^{28}}{67108864\pi^{7}}.$
7 Berndt-type Integrals via Barnes Multiple Zeta Functions
In this section, we evaluate Barnes multiple zeta function through the results of Berndt-type integrals. Building upon the work of Bradshaw and Vignat [7], who evaluated generalized Berndt-type integrals (1.1) using the (alternating) Barnes multiple zeta function, we present key results. For instance, they established ([7, Prop. 2]):
$\displaystyle\int_{0}^{\infty}\displaystyle\frac{x^{a}dx}{(\cos x-\cosh x)^{b}}=2^{b}\Gamma(a+1)\zeta_{2b}(a+1,b|(1+i,1-i)^{b})\quad(a\geq 2b,b\geq 1)$ (7.89)
and
$\displaystyle\int_{0}^{\infty}\displaystyle\frac{x^{a}dx}{(\cos x+\cosh x)^{b}}=2^{b}\Gamma(a+1){\bar{\zeta}}_{2b}(a+1,b|(1+i,1-i)^{b})\quad(a\geq 0,b\geq 1),$ (7.90)
[…]
$\zeta_{N}(s,\omega|{\bf a}_{N}):=\displaystyle\sum_{n_{1}\geq 0,\ldots,n_{N}\geq 0}\displaystyle\frac{1}{(\omega+n_{1}a_{1}+\cdots+n_{N}a_{N})^{s}}\quad(\Re(\omega)>0,\Re(s)>N)$ (7.91)
and
${\bar{\zeta}}_{N}(s,\omega|{\bf a}_{N}):=\displaystyle\sum_{n_{1}\geq 0,\ldots,n_{N}\geq 0}\displaystyle\frac{(-1)^{n_{1}+\cdots+n_{N}}}{(\omega+n_{1}a_{1}+\cdots+n_{N}a_{N})^{s}}\quad(\Re(\omega)>0,\Re(s)>N-1),$ (7.92)
respectively.
Bradshaw and Vignat [7] provided integral representations for these functions, recovering a known result [15, Eq. (3.2)].
Proposition 7.1.
Let $\Re(s)>N,\Re(w)>0$, and $\Re(a_{j})>0$ for $j=1,\ldots,N$. Then
$\zeta_{N}(s,w|a_{1},\ldots,a_{N})=\displaystyle\frac{1}{\Gamma(s)}\displaystyle\int_{0}^{\infty}u^{s-1}e^{-wu}\prod_{j=1}^{N}(1-e^{-a_{j}u})^{-1}du.$
The alternating version satisfies
$\widetilde{\zeta}_{N}(s,w|a_{1},\ldots,a_{N})=\displaystyle\frac{1}{\Gamma(s)}\displaystyle\int_{0}^{\infty}u^{s-1}e^{-wu}\prod_{j=1}^{N}(1+e^{-a_{j}u})^{-1}du.$
Theorem 7.2.
For positive integer $m$, we have
$\Gamma\left(4m\right)\zeta_{4}\left(4m,3|\mathbf{c}_{4};\boldsymbol{\sigma}_{4}\right)\in\mathbb{Q}\displaystyle\frac{\Gamma^{8m-4}}{\pi^{2m-1}}+\displaystyle\frac{\mathbb{Q}}{\sqrt{2}}\displaystyle\frac{\Gamma^{8m-2}}{\pi^{(4m-1)/2}}+\mathbb{Q}\displaystyle\frac{\Gamma^{8m}}{\pi^{2m}}+\displaystyle\frac{\mathbb{Q}}{\sqrt{2}}\displaystyle\frac{\Gamma^{8m+2}}{\pi^{(4m+3)/2}}+\mathbb{Q}\displaystyle\frac{\Gamma^{8m+4}}{\pi^{2m+3}},$
where $\mathbf{c}_{4}=\left(2+2i,2-2i,1+i,1-i\right)$ and $\boldsymbol{\sigma}_{4}=\left(\left\{1\right\}^{2},\left\{-1\right\}^{2}\right)$.
Proof.
Following [7, Prop. 2], we obtain
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{s-1}e^{-\omega x}}{\prod_{i=1}^{M}{\sinh\left(a_{i}x\right)}\prod_{j=1}^{N}{\cosh\left(b_{j}x\right)}}}\mathrm{d}x$ $=\displaystyle\sum_{\tiny\begin{array}[]{c}n_{1},\ldots,n_{M}\geq 0\\ k_{1},\ldots,k_{N}\geq 0\\ \end{array}}{\displaystyle\frac{2^{M+N}\Gamma\left(s\right)\left(-1\right)^{k_{1}+\cdots+k_{N}}}{\left(w+\displaystyle\sum_{i=1}^{M}{a_{i}+\displaystyle\sum_{i=1}^{N}{b_{i}+}2\left(\displaystyle\sum_{i=1}^{M}{a_{i}n_{i}}+\displaystyle\sum_{i=1}^{N}{b_{i}k_{i}}\right)}\right)^{s}}}$
$=2^{M+N}\Gamma\left(s\right)\zeta_{M+N}\left(s,w+\displaystyle\sum_{i=1}^{M}{a_{i}+}\displaystyle\sum_{i=1}^{N}{b_{i}}|\boldsymbol{c}_{M+N},\boldsymbol{\sigma}_{M+N}\right),$
where $\boldsymbol{c}_{M+N}=\left(2a_{1},\ldots,2a_{M},2b_{1},\ldots,2b_{N}\right)$ and $\boldsymbol{\sigma}_{M+N}=\left(\left\{1\right\}^{M},\left\{-1\right\}^{N}\right).$
Applying this to our integral
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m-1}\mathrm{d}x}{\left[\cosh(2x)-\cos(2x)\right]\left[\cosh x+\cos x\right]}}$
$=\displaystyle\frac{1}{4}\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m-1}\mathrm{d}x}{\sinh\left[\left(1+i\right)x\right]\sinh\left[\left(1-i\right)x\right]\cosh\left(\displaystyle\frac{1+i}{2}x\right)\cosh\left(\displaystyle\frac{1-i}{2}x\right)}}$
$=4\Gamma\left(4m\right)\zeta_{4}\left(4m,3|\mathbf{c}_{4},\boldsymbol{\sigma}_{4}\right),$
where $\mathbf{c}_{4}=\left(2+2i,2-2i,1+i,1-i\right)$ and $\boldsymbol{\sigma}_{4}=\left(\left\{1\right\}^{2},\left\{-1\right\}^{2}\right)$.
Finally, using Theorem 6.1 yields the desired evaluation. ∎
Employing the computational framework of Mathematica in conjunction with Theorem 7.2, we perform explicit calculations for integer values $m=1,2,3,4$.
Example 7.3.
Let $\Gamma=\Gamma(1/4)$. Then
$\zeta_{4}\left(4,3|\mathbf{c}_{4},\boldsymbol{\sigma}_{4}\right)$ $=\displaystyle\frac{\Gamma^{12}}{393216\pi^{5}}-\displaystyle\frac{\Gamma^{10}}{98304\sqrt{2}\pi^{7/2}}+\displaystyle\frac{\Gamma^{8}}{49152\pi^{2}}-\displaystyle\frac{\Gamma^{6}}{4096\sqrt{2}\pi^{3/2}}+\displaystyle\frac{\Gamma^{4}}{2048\pi},$
$\zeta_{4}\left(8,3|\mathbf{c}_{4},\boldsymbol{\sigma}_{4}\right)$ $=\displaystyle\frac{13\Gamma^{20}}{21139292160\pi^{7}}-\displaystyle\frac{29\Gamma^{18}}{7046430720\sqrt{2}\pi^{11/2}}+\displaystyle\frac{\Gamma^{16}}{146800640\pi^{4}}-\displaystyle\frac{3\Gamma^{14}}{41943040\sqrt{2}\pi^{7/2}}$
$\quad+\displaystyle\frac{\Gamma^{12}}{5242880\pi^{3}},$
$\zeta_{4}\left(12,3|\mathbf{c}_{4},\boldsymbol{\sigma}_{4}\right)$ $=\displaystyle\frac{241\Gamma^{28}}{1190564934451200\pi^{9}}-\displaystyle\frac{2147\Gamma^{26}}{1587419912601600\sqrt{2}\pi^{15/2}}+\displaystyle\frac{\Gamma^{24}}{442918502400\pi^{6}}$
$\quad-\displaystyle\frac{61\Gamma^{22}}{2576980377600\sqrt{2}\pi^{11/2}}+\displaystyle\frac{\Gamma^{20}}{16106127360\pi^{5}},$
$\zeta_{4}\left(16,3|\mathbf{c}_{4},\boldsymbol{\sigma}_{4}\right)$ $=\displaystyle\frac{55711\Gamma^{36}}{832062021389254656000\pi^{11}}-\displaystyle\frac{40487\Gamma^{34}}{90564573756653568000\sqrt{2}\pi^{19/2}}$
$\quad+\displaystyle\frac{\Gamma^{32}}{1340029796352000\pi^{8}}-\displaystyle\frac{179\Gamma^{30}}{22869841857740800\sqrt{2}\pi^{15/2}}+\displaystyle\frac{11\Gamma^{28}}{536011918540800\pi^{7}},$
where $\mathbf{c}_{4}=\left(2+2i,2-2i,1+i,1-i\right)$ and $\boldsymbol{\sigma}_{4}=\left(\left\{1\right\}^{2},\left\{-1\right\}^{2}\right)$.
Conflict of Interest. The authors declare no conflict of interest regarding the publication of this article.
Data Availability No new data were generated or analyzed in this study, and therefore data sharing is not applicable.
Use of AI Tools Declaration. The authors confirm that no artificial intelligence (AI) tools were used in the creation of this work.